OpenStudy (homeworksucks):

A net force of 5.75 x10^3 N [180 degrees] acts on a mass of 23.0 kg. If, while the force acts, the mass travels through a displacement of 360 m[210 degrees]. what work did the net force do on the object.

Parth (parthkohli):

\[{\bf W = F \times s = |F||s|}\sin\theta\]